OpenStudy (thedj4jc):

This question has me completely lost. The sum of the digits of a two-digit number is 12. The number formed by interchanging the digits is 54 more than the original number. What is the original number? 39 48 57

OpenStudy (anonymous):

t + u = 12 10u + t = 54 + 10t + u 9u - 9t = 54 u - t = 6 u + t = 12 t = 12 - u u - t = 6 u - (12 - u) = 6 u - 12 + u = 6 2u = 6 + 12 2u = 18 u = 9 9 + t = 12 t = 12 - 9 t = 3

OpenStudy (anonymous):

original: 39 reversal: 93 The number formed by interchanging the digits is 54 more than the original number 39 + 54 = 93

OpenStudy (thedj4jc):

Thank you!!!!

OpenStudy (anonymous):

your welcome! did you understand how i got the answer?

OpenStudy (thedj4jc):

Kinda